# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v2

@@ 1 Introduction
where $D$ denotes a suitable distance on the space of probability measures. In this framework $G_{\theta}$ is a conditional generative model, which in particular also depends on $y$. For instance, this is done in the framework of conditional (stochastic) normalizing flows [7, 28, 27, 59], conditional GANs [44] or conditional gradient flows for the Wasserstein metric [20, 26]. In practice however, we mostly look for the posteriors for single points $\tilde{y}$. Recently in [3], it was shown that the posterior error for single points $\tilde{y}$ goes to zero if the expected error to the posterior $\mathbb{E}_{Y}\left[W_{1}(P_{X|Y=y},P_{Z|Y=y})\right]$ is small in [3], where $W_{1}$ denotes the Wasserstein distance [56]. This shows that it is important to investigate the relation between $\mathbb{E}_{Y}\left[W_{1}(P_{X|Y=y},P_{Z|Y=y})\right]$ and $W_{1}(P_{Y,Z},P_{Y,X})$.
In [35], the authors investigated the relation between the distance of the joint measures $D(P_{Y,Z},P_{Y,X})$ and the expected error of the posteriors $\mathbb{E}_{Y}\left[D(P_{Z|Y=y},P_{X|Y=y})\right].$ For the Kullback–Leibler divergence $D=\text{KL}$, it follows by the chain rule [18, Theorem 2.5.3] that
@@
holds true, in contrast to the equality claim in [35, Theorem 2]. A simple counterexample is given in Appendix A. Intuitively, strict inequality can arise when the optimal transport (OT) plan needs to transport mass in the $Y$-component. This is the motivation for considering only plans that do not have mass transport in the $Y$-component. This leads us to the definition of conditional Wasserstein distances $W_{p,Y}$, where admissible transport plans are restricted to the set $\Gamma_{Y}^{4}=\Gamma_{Y}^{4}(P_{Y,X},P_{Y,Z})$ of 4-plans $\alpha$ fulfilling in addition $(\pi^{1,3})_{\sharp}\alpha=\Delta_{\sharp}P_{Y}$, where $\Delta(y)=(y,y)$ is the diagonal map:
@@
Inspired by [35], we show that this conditional Wasserstein distance indeed fulfills
@@
Further, we prove results on geodesics similar as in [6] for our conditional Wasserstein distance: we show the connections to the continuity equation, verify that there exists a velocity field with no mass transport in $Y$-direction and recover a corresponding ODE formulation. Indeed, this conditional Wasserstein distance can be used to explain a numerical observation made by [20, 26], namely that rescaling the $Y$-component leads to velocity fields with no mass transport in Y-direction in the limit. Using these ideas, we propose to relax the conditional Wasserstein distance to allow "small amounts" of mass transport in $Y$-direction.
Then, we use our insights to design efficient posterior sampling algorithms. By leveraging recent ideas of flow matching, see [39, 40], we design Bayesian OT flow matching. Note that the recent approaches of [60, 58] do not respect the OT in $X$-direction as they always choose a random coupling between $x$ and $z$ for each observation $y$ coupling. This leads to unfortunate situations, where the optimal $Y$-diagonal coupling is not recovered even between Gaussians when we approximate them by empirical measures, see Example 9. We use our proposed Bayesian OT flow matching and verify its advantages on a Gaussian mixture toy problem and on class conditional image generation on the CIFAR10 dataset.
@@
  • We show that the conditional Wasserstein distance can be used in conditional generative approaches and demonstrate the advantages on MNIST particle flows [26, 4]. We propose a version of OT flow matching [55, 49] for inverse problems which uses a relaxed version of our conditional Wasserstein distance, and show that it overcomes obstacles, explained in Example 9, from previous flow matching versions for inverse problems [58].
Related work
Our work is in the intersection of conditional generative modelling [1, 7, 44] and (computational) OT [48, 56]. The recent work [24, Theorem 2] derives an inequality based on restricting the admissible couplings in the their OT formulation to so-called conditional sub-couplings. Note that their reformulation is only a restatement of the expected value, but does not relate it to the joint distributions. Those authors also look for geodesics in the Wasserstein space, but pursue a different approach. While we relate it to the velocity fields in gradient flows, they pursue an autoencoder/GAN idea.
@@
The closest work, which appeared after our first version of this paper, is [31]. Here the authors define the conditional optimal transport problem and calculate its dual. Their work is more focused on the infinite dimensional setting, whereas we consider the velocity fields needed for the flow matching application. There are two other concurrent works that treat objects similar to what is presented in this paper. [10] study gradient flows for conditional Wasserstein spaces in the case where $P_{Y}$ is the Lesbesgue measure on $[0,1]$ in order to analyse the training of infinitely deep and wide ResNets. Another preprint appearing after the ArXiv submission of the present paper, which also uses conditional OT for flow matching, is [34]. They come up with a similar loss function, which also uses that the velocity field should not transport mass in the Y-component. Another concurrent work is [33], where they use a generalized continuity equation to extend the flow matching framework to the matrix valued case, where they use it for style transfer. Theoretically our paper is more focused on finding geodesics where no mass in $Y$ is transported whereas they look for translations between classes.
@@
In the OT literature, there has been a collection of class conditional OT distances used in domain adaption [45, 50]. In particular, conditional OT as in [53] is relevant as they consider OT plans for each condition $y$ minimizing $\mathbb{E}_{y}[W_{1}(P_{X|Y=y},G(\cdot,y)_{\#}P_{Z})]$. However they relax their problem using a KL divergence. The works on Wasserstein gradient flows [6, 25] investigate conditional Wasserstein distances from a different point of view for defining the so-called geometric tangent space of the 2-Wasserstein space. Geometric tangent spaces play a crucial role in Wasserstein flows of maximum mean discrepancies with Riesz kernels in [29] and their neural variants in [4]. In [26, Remark 7], an inequality between the joint Wasserstein and the expected value over the conditionals is derived, where the result requires compactly supported measures and certain regularity of the associated posterior densities. In [14], the supervised training of conditional Monge maps is proposed, for which the authors solved the dual problem using convex neural networks. The authors of [41] also consider an amortized objective between the conditional distributions and propose a relaxation, which only needs samples from the joint distribution involving maximum mean discrepancies. Numerically, we first verify our theoretical statements based on particle flows, which were also used in [4, 26]. Further, we apply our framework to solve inverse problems using Bayesian flow matching [58, 60] and OT flow matching [39, 40, 55, 49].
@@
This paper is an extension of our first ArXiv version [15] on conditional Wasserstein distances with more focus on the continuity equation and flow equation for geodesics as well as flow matching.
Outline of the paper
@@
In Section 2, we recall preliminaries from OT. Then, in Section 3, we introduce conditional Wasserstein distances of joint probability measures, and show their relation to the expectation over the Wasserstein distance of the conditional probabilities. Moreover, we highlight the connection to work on geometric tangent spaces. In Section 4, we calculate the dual of our conditional Wasserstein-1 distance and show how a loss function used in the conditional Wasserstein GAN literature arises in a natural way. In Section 5, we deal with geodesics with respect to the conditional Wasserstein distance, prove properties of the corresponding velocity fields, showing that they vanish in the $Y$-component, and show existence for flow ODEs. We propose a relaxation of the conditional Wasserstein distance which appears to be useful for numerical computations in Section 6. We combine our findings with OT flow matching to get Bayesian flow matching in Section 7. Finally, in Section 8, we present numerical results: we verify a convergence result for an approximation of the conditional Wasserstein distance using particle flows to MNIST, and demonstrate the advantages of our Bayesian OT flow matching procedure on a Gaussian mixture model toy example and on CIFAR10 class-conditional image generation. All proofs are postponed to the appendices.
@@ 2 Preliminaries
Throughout this paper, we will use the following notation. These are basics from from optimal transport theory and can be found in [56]. By $\mathcal{P}(\mathbb{X})$, we denote the set of probability measures on $\mathbb{X}\subseteq\mathbb{R}^{n}$ and by $\mathcal{P}_{p}(\mathbb{X})$, $p\in[1,\infty)$ the subset of measures with finite $p$-th moments. For $\mu\in\mathcal{P}(\mathbb{X})$ and a measurable function $F:\mathbb{X}\to\mathbb{Y}$, we define the the push forward measure by $F_{\sharp}\mu=F\circ\mu^{-1}$. For a product space $\prod_{i=1}^{K}\mathbb{X}_{i}$, we denote the projection onto the $i_{1},\ldots,i_{k}$-th component by $\pi^{i_{1},\ldots,i_{k}}$. The Wasserstein-$p$ metric [56] on $\mathcal{P}_{p}(\mathbb{X})$ is given by
@@
[…]
@@
Further, this optimal map is related to the optimal transport plan $\gamma$ in (3) by $\gamma=(\mathrm{Id},T)_{\sharp}\mu,$ see [56]. The same holds true for empirical measures with the same number of points, see [48, Proposition 2.1]. In this paper, we ask for relations between joint and posterior probabilities: for random variables $X,Z\in B\subseteq\mathbb{R}^{m}$ and $Y\in A\subseteq\mathbb{R}^{d}$, we are interested in Wasserstein distances between $P_{Y,X},P_{Y,Z}\in\mathcal{P}_{p}(A\times B)$ and $P_{X|Y=y},P_{Z|Y=y}\in\mathcal{P}_{p}(B)$. Since $\pi^{1}_{\sharp}P_{Y,X}=P_{Y}$ as well as $\pi^{1}_{\sharp}P_{Y,Z}=P_{Y}$, we see that the joint probabilities belong indeed to the subset
@@
For $p=2$, this set was considered as set of velocity plans at $P_{Y}$ in [6, Sect. 12.4] and [25, Sect. 4]. It is the basis for defining the so-called geometric tangent space of $\mathcal{P}_{2}(\mathbb{R}^{d})$ which was used, e.g. in [29, 4] for handling (neural) Wasserstein gradient flows of maximum mean discrepancies.
We will frequently apply the disintegration formula [6, Theorem 5.3.1] which says that for a measure $\gamma\in\mathcal{P}(A\times B)$ with $\pi^{1}_{\sharp}\gamma=\mu\in\mathcal{P}(A)$, there exists a $\mu$-a.e. uniquely determined Borel family of probability measures $(\gamma_{y})_{y\in A}$ such that
@@ 3 Conditional Wasserstein distance
Note that the definition of $\alpha$ in $iii)$ already appears in the proof of [35, Theorem 2]. For $p=2$, it was shown in [6, Sect. 12.4] and [25, Sect. 4] that the square root of the right-hand side in (10) is a metric on $\mathcal{P}_{2,Y}(A\times B)$. The proof can be generalized in a straightforward way for $p\in[1,\infty)$. Thus, by Proposition 1, we have the following corollary.
Corollary 2.
@@
Interestingly, for $p=2$, there was also given an equivalent definition by [25] of $\mathcal{W}_{p,Y}$, namely
@@
The relation between the admissible 3-plans and 4-plans is given in the following proposition for which a proof can be found in the appendix.
Proposition 3.
@@ 4 Dual formulation of $W_{1,Y}$ and relation to GAN loss
Interestingly, the conditional Wasserstein-1 distance recovers loss functions in the conditional Wasserstein GAN literature [1, 35, 42]. Wasserstein GANs [8] aim to sample from a target distribution $P_{X}$ based on a simpler distribution $P_{Z}$. A generator $G=G_{\theta}$ is learned such that the Wasserstein-1 distance in its dual formulation
@@
Usually, the the 1-Lipschitz condition is enforced via so-called weight-clipping [8].
In [1], this approach was generalized to inverse problems. Assume that $A\subset\mathbb{R}^{d}$ and $B\subset\mathbb{R}^{m}$ are compact sets throughout this section. For given $y\in A$, let $h(y,\cdot)\in\text{Lip}_{1}$ be a minimizer in
@@
where the maximum is taken over functions $h$ which are Lipschitz-1 continuous in the second variable. However, exchanging maximum and expectation value requires that $(y,x)\mapsto h(y,x)$ is measurable which is not immediate. This ,,gap” was fixed under stronger assumptions, e.g. on the posterior, in [42].
In this section, we show that the dual formulation of the conditional Wasserstein distance $W_{1,Y}$ leads naturally to the desired loss on the right-hand side of (13) for an appropriate regular function set for $h$. More precisely, we have the following theorem which is proved in Appendix C.
Theorem 4.

### v3

@@ Abstract
Keywords: Conditional Wasserstein distance, posterior sampling, flow matching, inverse problems, generative modelling
@@ 1 Introduction
where $D$ denotes a suitable distance on the space of probability measures. In this framework $G_{\theta}$ is a conditional generative model, which in particular also depends on $y$. For instance, this is done in the framework of conditional (stochastic) normalizing flows (Ardizzone et al., 2019; Hagemann et al., 2022b, a; Winkler et al., 2019), conditional GANs (Mirza and Osindero, 2014) or conditional gradient flows for the Wasserstein metric (Du et al., 2023; Hagemann et al., 2024). In practice however, we mostly look for the posteriors for single points $\tilde{y}$. Recently in (Altekrüger et al., 2023), it was shown that the posterior error for single points $\tilde{y}$ goes to zero if the expected error to the posterior $\mathbb{E}_{Y}\left[W_{1}(P_{X|Y=y},P_{Z|Y=y})\right]$ is small in (Altekrüger et al., 2023), where $W_{1}$ denotes the Wasserstein distance (Villani, 2009). This shows that it is important to investigate the relation between $\mathbb{E}_{Y}\left[W_{1}(P_{X|Y=y},P_{Z|Y=y})\right]$ and $W_{1}(P_{Y,Z},P_{Y,X})$.
In (Kim et al., 2023), the authors investigated the relation between the distance of the joint measures $D(P_{Y,Z},P_{Y,X})$ and the expected error of the posteriors $\mathbb{E}_{Y}\left[D(P_{Z|Y=y},P_{X|Y=y})\right].$ For the Kullback–Leibler divergence $D=\text{KL}$, it follows by the chain rule (Cover, 2005, Theorem 2.5.3) that
@@
holds true, in contrast to the equality claim in (Kim et al., 2023, Theorem 2). A simple counterexample is given in Appendix A. Intuitively, strict inequality can arise when the optimal transport (OT) plan needs to transport mass in the $Y$-component. This is the motivation for considering only plans that do not have mass transport in the $Y$-component. This leads us to the definition of conditional Wasserstein distances $W_{p,Y}$, where admissible transport plans are restricted to the set $\Gamma_{Y}^{4}=\Gamma_{Y}^{4}(P_{Y,X},P_{Y,Z})$ of 4-plans $\alpha$ fulfilling in addition $(\pi^{1,3})_{\sharp}\alpha=\Delta_{\sharp}P_{Y}$, where $\Delta(y)=(y,y)$ is the diagonal map:
@@
Inspired by (Kim et al., 2023), we show that this conditional Wasserstein distance indeed fulfills
@@
Further, we prove results on geodesics similar as in (Ambrosio et al., 2005) for the conditional Wasserstein distance: we show the connections to the continuity equation, verify that there exists a velocity field with no mass transport in $Y$-direction and recover a corresponding ODE formulation. Indeed, this conditional Wasserstein distance can be used to explain a numerical observation made by (Du et al., 2023; Hagemann et al., 2024), namely that rescaling the $Y$-component leads to velocity fields with no mass transport in Y-direction in the limit. Using these ideas, we propose to relax the conditional Wasserstein distance to allow "small amounts" of mass transport in $Y$-direction.
Then, we use our insights to design efficient posterior sampling algorithms. By leveraging recent ideas of flow matching, see (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023; Lipman et al., 2023), we design Bayesian OT flow matching. Note that the recent approaches of (Zheng et al., 2023; Wildberger et al., 2023) do not respect the OT in $X$-direction as they always choose a random coupling between $x$ and $z$ for each observation $y$. This leads to unfortunate situations, where the optimal $Y$-diagonal coupling is not recovered even between Gaussians when we approximate them by empirical measures, see Example 1. We use our proposed Bayesian OT flow matching and verify its advantages on a Gaussian mixture toy problem and on class conditional image generation on the CIFAR10 dataset.
@@
  • We show that the conditional Wasserstein distance can be used in conditional generative approaches and demonstrate the advantages on MNIST particle flows (Hagemann et al., 2024; Altekrüger et al., 2023). We propose a version of OT flow matching (Tong et al., 2023; Pooladian et al., 2023) for inverse problems which uses a relaxed version of our conditional Wasserstein distance, and show that it overcomes obstacles, explained in Example 1, from previous flow matching versions for inverse problems (Wildberger et al., 2023; Albergo et al., 2024).
@@
Related work Our work is in the intersection of conditional generative modeling (Adler and Öktem, 2018; Ardizzone et al., 2019; Mirza and Osindero, 2014) and (computational) OT (Peyré and Cuturi, 2019; Villani, 2009). The recent work (Kim et al., 2023, Theorem 2) derives an inequality based on restricting the admissible couplings in the their OT formulation to so-called conditional sub-couplings. Note that their reformulation is only a restatement of the expected value, but does not relate it to the joint distributions. Those authors also look for geodesics in the Wasserstein space, but pursue a different approach. While we relate it to the velocity fields in gradient flows, they pursue an autoencoder/GAN idea. To the best of our knowledge, the first work which considered conditional Wasserstein distances was Kloeckner (2021). Related absolutely continuous curves were discussed in Peszek and Poyato (2023) including the existence of vector fields for absolutely continuous curves. Peszek and Poyato (2023) were mainly interested in absolutely continuous curves stemming from gradient flows and not in geodesics. The closest work, which appeared after our first version of this paper, is (Hosseini et al., 2024). Here the authors define the conditional optimal transport problem and calculate its dual. Their work is more focused on the infinite dimensional setting, whereas we consider the velocity fields needed for the flow matching application. There are two other concurrent works that treat objects similar to what is presented in this paper. (Barboni et al., 2024) study gradient flows for conditional Wasserstein spaces in the case where $P_{Y}$ is the Lesbesgue measure on $[0,1]$ in order to analyse the training of infinitely deep and wide ResNets. Another preprint appearing after the ArXiv submission of the present paper, which also uses conditional OT for flow matching, is (Kerrigan et al., 2024). They come up with a similar loss function, which also uses that the velocity field should not transport mass in the Y-component. Another concurrent work is (Isobe et al., 2024), where they use a generalized continuity equation to extend the flow matching framework to the matrix valued case, where they use it for style transfer. Theoretically our paper is more focused on finding geodesics where no mass in $Y$ is transported whereas they look for translations between classes.
@@
In the OT literature, there has been a collection of class conditional OT distances used in domain adaption (Nguyen et al., 2022; Rakotomamonjy et al., 2022). In particular, conditional OT as in (Tabak et al., 2021) is relevant as they consider OT plans for each condition $y$ minimizing $\mathbb{E}_{y}[W_{1}(P_{X|Y=y},G(\cdot,y)_{\#}P_{Z})]$. However they relax their problem using a KL divergence. The works on Wasserstein gradient flows (Ambrosio et al., 2005; Gigli, 2008) investigate conditional Wasserstein distances from a different point of view for defining the so-called geometric tangent space of the 2-Wasserstein space. Geometric tangent spaces play a crucial role in Wasserstein gradient flows of maximum mean discrepancies with Riesz kernels in (Hertrich et al., 2023) and their neural variants in (Altekrüger et al., 2023). In (Hagemann et al., 2024, Remark 7), an inequality between the joint Wasserstein and the expected value over the conditionals is derived, where the result requires compactly supported measures and certain regularity of the associated posterior densities. In (Bunne et al., 2022), the supervised training of conditional Monge maps is proposed, for which the authors solved the dual problem using convex neural networks. The authors of (Manupriya et al., 2023) also consider an amortized objective between the conditional distributions and propose a relaxation, which only needs samples from the joint distribution involving maximum mean discrepancies. Numerically, we first verify our theoretical statements based on particle flows, which were also used in (Altekrüger et al., 2023; Hagemann et al., 2024). Further, we apply our framework to solve inverse problems using Bayesian flow matching (Wildberger et al., 2023; Zheng et al., 2023) and OT flow matching (Albergo et al., 2024; Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023; Pooladian et al., 2023).
@@
This paper is an extension of our first ArXiv version (Chemseddine et al., 2023) on conditional Wasserstein distances with more focus on the continuity equation and flow equation for geodesics as well as flow matching.
@@
Outline of the paper In Section 2, we recall preliminaries from OT. Then, in Section 3, we introduce conditional Wasserstein distances of joint probability measures, and show their relation to the expectation over the Wasserstein distance of the conditional probabilities. Moreover, we highlight the connection to work on geometric tangent spaces. In Section 4, we calculate the dual of our conditional Wasserstein-1 distance and show how a loss function used in the conditional Wasserstein GAN literature arises in a natural way. In Section 5, we deal with geodesics with respect to the conditional Wasserstein distance, prove properties of the corresponding velocity fields, showing that they vanish in the $Y$-component, and show existence for flow ODEs. We propose a relaxation of the conditional Wasserstein distance which appears to be useful for numerical computations in Section 6. We combine our findings with OT flow matching to get Bayesian flow matching in Section 7. Finally, in Section 8, we present numerical results: we verify a convergence result for an approximation of the conditional Wasserstein distance using particle flows to MNIST, and demonstrate the advantages of our Bayesian OT flow matching procedure on a Gaussian mixture model toy example and on CIFAR10 class-conditional image generation. All proofs are postponed to the appendices.
@@ 2 Preliminaries
Throughout this paper, we will use the following notation. These are basics from from optimal transport theory and can be found in (Villani, 2009). By $\mathcal{P}(\mathbb{X})$, we denote the set of probability measures on $\mathbb{X}\subseteq\mathbb{R}^{n}$ and by $\mathcal{P}_{p}(\mathbb{X})$, $p\in[1,\infty)$ the subset of measures with finite $p$-th moments. For $\mu\in\mathcal{P}(\mathbb{X})$ and a measurable function $F:\mathbb{X}\to\mathbb{Y}$, we define the push forward measure by $F_{\sharp}\mu=F\circ\mu^{-1}$. For a product space $\prod_{i=1}^{K}\mathbb{X}_{i}$, we denote the projection onto the $i_{1},\ldots,i_{k}$-th component by $\pi^{i_{1},\ldots,i_{k}}$. The Wasserstein-$p$ metric (Villani, 2009) on $\mathcal{P}_{p}(\mathbb{X})$ is given by
@@
where $\Gamma=\Gamma(\mu,\nu)$ denotes the set of all probability measures $\gamma\in\mathcal{P}(\mathbb{X}\times\mathbb{X})$ with marginals $\pi^{1}_{\sharp}\gamma=\mu$ and $\pi^{2}_{\sharp}\gamma=\nu$ and $\|\cdot\|$ is the Euclidean distance on $\mathbb{R}^{n}$, see (Villani, 2009). If $\mu\in\mathcal{P}_{p}(\mathbb{X})$ is absolutely continuous, then, for $p\in(1,\infty)$, there exists a unique optimal transport map $T\in L^{p}_{\mu}(\mathbb{X},\mathbb{X})$, also known as Monge map, which solves
@@
Further, this optimal map is related to the optimal transport plan $\gamma$ in (3) by $\gamma=(\mathrm{Id},T)_{\sharp}\mu,$ see (Villani, 2009). The same holds true for empirical measures with the same number of points, see (Peyré and Cuturi, 2019, Proposition 2.1). In this paper, we ask for relations between joint and posterior probabilities: for random variables $X,Z\in B\subseteq\mathbb{R}^{m}$ and $Y\in A\subseteq\mathbb{R}^{d}$, we are interested in Wasserstein distances between $P_{Y,X},P_{Y,Z}\in\mathcal{P}_{p}(A\times B)$ and $P_{X|Y=y},P_{Z|Y=y}\in\mathcal{P}_{p}(B)$. Since $\pi^{1}_{\sharp}P_{Y,X}=P_{Y}$ as well as $\pi^{1}_{\sharp}P_{Y,Z}=P_{Y}$, we see that the joint probabilities belong indeed to the subset
@@
For $p=2$, this set was considered as set of velocity plans at $P_{Y}$ in (Ambrosio et al., 2005, Sect. 12.4) and (Gigli, 2008, Sect. 4). It is the basis for defining the so-called geometric tangent space of $\mathcal{P}_{2}(\mathbb{R}^{d})$ which was used, e.g. in (Hertrich et al., 2023; Altekrüger et al., 2023) for handling (neural) Wasserstein gradient flows of maximum mean discrepancies.
We will frequently apply the disintegration formula (Ambrosio et al., 2005, Theorem 5.3.1) which says that for a measure $\gamma\in\mathcal{P}(A\times B)$ with $\pi^{1}_{\sharp}\gamma=\mu\in\mathcal{P}(A)$, there exists a $\mu$-a.e. uniquely determined Borel family of probability measures $(\gamma_{y})_{y\in A}$ such that
@@ 3 Conditional Wasserstein Distance
Note that the definition of $\alpha$ in $iii)$ already appears in the proof of (Kim et al., 2023, Theorem 2). For $p=2$, it was shown in (Ambrosio et al., 2005, Sect. 12.4) and (Gigli, 2008, Sect. 4) that the square root of the right-hand side in (10) is a metric on $\mathcal{P}_{2,Y}(A\times B)$. The proof can be generalized in a straightforward way for $p\in[1,\infty)$. Thus, by Proposition 1, we have the following corollary.
Corollary 2
@@
Interestingly, for $p=2$, there was also given an equivalent definition by (Gigli, 2008) of $\mathcal{W}_{p,Y}$, namely
@@
The relation between the admissible 3-plans and 4-plans is given in the following proposition, for which a proof can be found in the appendix.
Proposition 3
@@ 4 Dual Formulation of $W_{1,Y}$ and Relation to GAN Loss
Interestingly, the conditional Wasserstein-1 distance recovers loss functions in the conditional Wasserstein GAN literature (Adler and Öktem, 2018; Kim et al., 2023; Martin, 2021). Wasserstein GANs (Arjovsky et al., 2017) aim to sample from a target distribution $P_{X}$ based on a simpler distribution $P_{Z}$. A generator $G=G_{\theta}$ is learned such that the Wasserstein-1 distance in its dual formulation
@@
Usually, the 1-Lipschitz condition is enforced via so-called weight-clipping (Arjovsky et al., 2017).
In (Adler and Öktem, 2018), this approach was generalized to inverse problems. Assume that $A\subset\mathbb{R}^{d}$ and $B\subset\mathbb{R}^{m}$ are compact sets throughout this section. For given $y\in A$, let $h(y,\cdot)\in\text{Lip}_{1}$ be a minimizer in
@@
where the maximum is taken over functions $h$ which are Lipschitz-1 continuous in the second variable. However, exchanging maximum and expectation value requires that $(y,x)\mapsto h(y,x)$ is measurable which is not immediate. This ,,gap” was fixed under stronger assumptions, e.g. on the posterior, in (Martin, 2021).
In this section, we show that the dual formulation of the conditional Wasserstein distance $W_{1,Y}$ leads naturally to the desired loss on the right-hand side of (13) for an appropriate regular function set for $h$. More precisely, we have the following theorem which is proved in Appendix C. Note that we can give the precise space $\mathcal{F}$, where the functions we take the supremum over belong to.
Theorem 4